OpenStudy (anonymous):

The amount of bacteria y in a petri dish grows according to the equation dy/dt=ky where k is a constant and t is measured in hours. If the ammount of bacteria triples in 10 hours, then k is apx....

OpenStudy (anonymous):

okay so lets back it up just a lil. we are soing to set the first eaquation with all the constants in place, expt d= 1

OpenStudy (anonymous):

\[\left(\begin{matrix}1Y \\ 1T\end{matrix}\right)=KY\]

OpenStudy (anonymous):

however the second amount of info says that y triples and T= 10 then we plug it in the equation. (3/10)=K*3

OpenStudy (anonymous):

inverse operations and yu get (3/10)(1/3)=K=.10